OpenStudy (shelby1290):

Find the coordinates of the centroid given the vertices of triangle PQR. P(-6,9), Q(6,1), R(-6,-7) Solution (-2,1)

OpenStudy (bibby):

the centroid of a triangle can be found using this formula: \((x,y)=(\dfrac{A_x+B_x+C_x}{3},\dfrac{A_y+B_y+C_y}{3})\) where \(P(-6,9)=(A_x,A_y)\)

OpenStudy (bibby):

basically, add up the x components and divide by 3 add the y's and divide by 3

OpenStudy (shelby1290):

I know that's one way of doing it but my teacher does it with a lot more steps!

OpenStudy (shelby1290):

@bibby She finds 2 midpoints. Then finds the slope and substitutes or eliminates

OpenStudy (bibby):

can you show me an example of how they do it? does it involve the distance formula

OpenStudy (bibby):

oh, gross

OpenStudy (shelby1290):

or something like that

OpenStudy (shelby1290):

yeaaa lol

OpenStudy (bibby):

can you give me a minute to google an example because that isn't really clear

OpenStudy (shelby1290):

should i draw out what I've done so far?

OpenStudy (bibby):

uh, that' works too

OpenStudy (bibby):

oh, so you find 2 midpoints draw the 2 medians to those points and then the point where they cross is the centroid

OpenStudy (shelby1290):